A 50g arrow moving at 20m/s strikes a pumpkin and drills a 35cm hole straight through it. When it emerges, the arrow has a speed of 8m/s. Assume uniform deceleration, find the time taken by the arrow to pass through the pumpkin.


1
Expert's answer
2022-03-09T15:55:05-0500

The time taken by the arrow to pass through the pumpkin

"t=\frac{2d}{v_i+v_f}\\
=\frac{2*0.35}{20+8}=0.025\:\rm s"